OpenStudy (anonymous):

how to factor 36a2x - b2

OpenStudy (ajprincess):

Is ur question \(36a^2x-b^2\)?

OpenStudy (anonymous):

no, 36a^2x-b^2

OpenStudy (ajprincess):

Are u sure there is x after a^2?

OpenStudy (anonymous):

yes

OpenStudy (ajprincess):

if 36a^2x-b^2 is ur question I dnt think it can be factorised.

OpenStudy (anonymous):

Thanks

OpenStudy (ajprincess):

yw. If ur question is 36a^2-b^2 then it can be factorised as follows. 36a^2-b^2 =6^2a^2-b^2 =(6a)^2-b^2 =(6a-b)(6a+b) a^2-b^2=(a-b)(a+b)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

now can you help me factor 16a^2 - 4b^2

OpenStudy (anonymous):

please

OpenStudy (ajprincess):

\(16a^2 - 4b^2\) \(=4(4a^2-b^2)\) \(=4(2^2a^2-b^2)\) \(=4((2a)^2-b^2)\) \(=4(2a-b)(2a+b)\)

OpenStudy (anonymous):

Thank you. Can you help me factor my last one. 36x^2y^6 - 16

OpenStudy (ajprincess):

Please follow the method I showed to factorise it.